2. If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
A. 2 hours and 24 minutes
B. 3 hours and 12 minutes
C. 3 hours and 44 minutes
D. 4 hours and 10 minutes
E. 4 hours and 33 minutes
Please show me how to solve this problem.

Answer by palanisamy(496)

Given, Sally can paint a house in 4 hours, and John can paint the same house in 6 hour.
Work done by Sally in one hour = 1/4
Work done by Jonh in one hour = 1/6
Total work done by both of them in one hour = (1/4+1/6)
= (3+2)/12
= 5/12
Time taken to paint the house when both of them paint together = 12/5 hours
=2 2/5 hours
=2h (2/5)*60 min
=2h 24 m